The ratio of daily wages of two workers is 4 : 3 and one gets daily $ 9 more than the other, what are their daily wages?
(a) $ 32 and $ 24 (b) $ 60 and $ 45 (c) $ 80 and $ 60 (d) $ 36 and $ 27

Answer by stanbon(75887)

The ratio of daily wages of two workers is 4 : 3 and one gets daily $ 9 more than the other, what are their daily wages?
-------------
Proportion:
4/3 = (x+9)/x
4x = 3x+27
x = 27
-----
x+9 = 36
-----------------------
Ans: d
Cheers,
Stan H.
===================
(a) $ 32 and $ 24 (b) $ 60 and $ 45 (c) $ 80 and $ 60 (d) $ 36 and $ 27